OpenStudy (anonymous):

Larry has taken out a loan for college. He started paying off the loan with a first payment of $150. Each month he pays, he wants to pay back 1.3 times as the amount he paid the month before. Explain to Larry how to represent his first 15 payments in sigma notation. Then explain how to find the sum of his first 15 payments, using complete sentences. Explain why this series is convergent or divergent.

OpenStudy (amistre64):

so whats your explanation so far?

OpenStudy (anonymous):

so far i have the equation a1-a1r^n/1-r which i think would accompany this question but I'm not sure

OpenStudy (anonymous):

so far, I'm just doing the math, trying to figure out what the formula may be

OpenStudy (amistre64):

that at least has something to do with a geometric series :)

OpenStudy (anonymous):

lol A for effort

OpenStudy (amistre64):

lets start out with a payment: P month1 = P month2 = (P) * k month3 = (P * k) * k month4 = ((P * k) * k) * k simplifing this stuff we see that month1 = P month2 = Pk month3 = Pk^2 month4 = Pk^3 .... month n = Pk^(n-1) agreed?

OpenStudy (anonymous):

YES, completely

OpenStudy (anonymous):

so now, all i do is substitute my values?

OpenStudy (amistre64):

well, then thats how the sequence is generates the sum of payments is just the sum from months 1 to n; and yes you would want to substitute your values into this generalization

OpenStudy (amistre64):

\[\sum_{n=1}^{j}Pk^{n-1}\] this of course can be written in several ways but is a good form as is

OpenStudy (amistre64):

since j=15, would you agree that the sum of 15 numbers is somehting that produces a fininte value?

OpenStudy (anonymous):

no of coarse not

OpenStudy (amistre64):

for example; spose you take out all the change from your piggy bank ... there may be hundreds of coins, even thousands .... or maybe just 2 coins but if we count the value of all those coins, we come to some knowable value right?

OpenStudy (anonymous):

yeah of coarse

OpenStudy (amistre64):

then we can count 15 payments and reach some known value ... it converges to some number

OpenStudy (amistre64):

you can use the formula you presented at first to determine that value, but in terms of what ive been using:\[S_j=P\frac{1-k^j}{1-k}\]

OpenStudy (anonymous):

exactly, it has to reach a certain point, it doest go on for infinity, right?

OpenStudy (amistre64):

right :)

OpenStudy (anonymous):

oh alright, i think i get it, so this is the sum? great!! :D

OpenStudy (amistre64):

with any luck :)

OpenStudy (anonymous):

so for 15 months, it would be 25092?

OpenStudy (amistre64):

thats correct

OpenStudy (amistre64):

i have to go pick up my kids, so have fun with this :)

OpenStudy (anonymous):

Thank yo so much!! Have a blessed day with your kids <3